OpenStudy (anonymous):

9x^2-12x+5=0 Solve quadratic equation by completing the square.

OpenStudy (anonymous):

i have an idea but its just not working

OpenStudy (wolf1728):

Move non-x term to the right 9x^2 -12x = -5 Divide by coefficient of x^2 x^2 -1.3333x = -(5/9) Get coefficient of x, divide by 2 then square it -1.3333 / 2 = -0.6666666667 (-0.6666666667)^2 = 0.4444444444 Add this to both sides x^2 -1.33333x + .44444 = -4.5555555556 mutiplying both sides by -1 -x^2 + 1.333333x = 4.555555 take square root of both sides

OpenStudy (anonymous):

My teacher says to keep in fraction form but this looks right thanks!

OpenStudy (wolf1728):

Okay - the answers by the way are: x = 1.6667 x = -0.33333 using quadratic formula

OpenStudy (anonymous):

thanks a lot!